# Knowledge Graph Large Language Model (KG-LLM) for Link Prediction

### v8

@@ Abstract
The task of multi-hop link prediction within knowledge graphs (KGs) stands as a challenge in the field of knowledge graph analysis, as it requires the model to reason through and understand all intermediate connections before making a prediction. In this paper, we introduce the Knowledge Graph Large Language Model (KG-LLM), a novel framework that leverages large language models (LLMs) for knowledge graph tasks. We first convert structured knowledge graph data into natural language and then use these natural language prompts to fine-tune LLMs to enhance multi-hop link prediction in KGs. By converting the KG to natural language prompts, our framework is designed to learn the latent representations of entities and their interrelations. To show the efficacy of the KG-LLM Framework, we fine-tune three leading LLMs within this framework, including Flan-T5, Llama2 and Gemma. Further, we explore the framework’s potential to provide LLMs with zero-shot capabilities for handling previously unseen prompts. Experimental results show that KG-LLM significantly improves the models’ generalization capabilities, leading to more accurate predictions in unfamiliar scenarios. Our code is available at https://anonymous.4open.science/r/KG-LLM-FED0.
@@ 1 Introduction
Historically, approaches to solving tasks related to KGs can trace their origins from embedding-based methods to more recent advancements with LLMs [28]. Initially, embedding-based methods played a crucial role, utilizing techniques to represent both entities and relations in a KG as low-dimensional vectors to address the link prediction task by preserving the structural and semantic integrity of the graph [2, 30, 5, 13]. As the field progressed, the integration of LLMs began to offer new paradigms, leveraging large amounts of data and advanced architectures to further enhance prediction capabilities and semantic understanding in KGs [1, 38, 39, 37, 22]. This transition shows a significant improvement from using purely mathematical representations to more context-aware methodologies that better understand the knowledge representation.
@@
To bridge these gaps, we propose the Knowledge Graph Large Language Model (KG-LLM), a novel approach to multi-hop link prediction. As illustrated in Figure 1, nodes in KGs are interconnected through specific relations. Initially, our framework takes input from the original KG dataset. After preprocessing, all paths in the KG will transform into chain-of-thought (CoT) prompts [33], each includes a series of relational statements, which can be represented as {Node 1 (has relation_$x$ with) Node 2, Node 2 (has relation_$y$ with) Node 3, etc.}. The complexity of the multi-hop problem is determined by its path length and the number of nodes. Through instruction fine-tuning (IFT) [31] of three Large Language Models (LLMs): Flan-T5 [31], Llama2 [24], and Gemma [23], our framework is ready to enhance multi-hop link prediction performance during the testing phase. Moreover, by integrating in-context learning (ICL) [35], the model not only improves but also has the capacity to tackle unseen prompts, showcasing our method’s innovativeness in addressing multi-hop link prediction challenges.
@@ 2 Related Work
Recently, researchers have used Graph Neural Network (GNN) models to solve various graph-related tasks, significantly advancing the field. Among different GNN models, Graph Attention Networks (GATs) have gained attention for their ability to weigh the importance of neighboring nodes, with models like wsGAT [6] demonstrating effectiveness in link prediction tasks. Additionally, Graph Convolutional Network (GCN)-based models have shown promising results; ConGLR [12] leverages context graphs and logical reasoning for improved inductive relation prediction, while ConvRot [11] integrates relational rotation and convolutional techniques to enhance link prediction performance in knowledge graphs (KGs). While the aforementioned approaches have achieved significant success, multi-hop link prediction remains an unsolved challenge.
Other than GNN models, recent development of large language models (LLMs), such as BERT [4], GPT [18], Llama [24], Gemini [23], and Flan-T5 [31] has also solved various KGs tasks, including link prediction. The text-to-text training approach makes LLMs particularly suitable for our generative multi-hop link prediction task. Recent studies, concurrent work such as GraphEdit [8], MuseGraph [21], and InstructGraph [27], have shown that natural language is effective for representing structural data for LLMs. Besides, training on large-scale data makes it possible for LLMs to generalize to unseen tasks or prompts that were not part of its training data [32].
Another advantage of LLM-based generative modeling is the Chain-of-Thought (CoT) reasoning ability [34]. It provides the flexibility of modifying the instruction, options, and exemplars to allow structured generation and prediction. The Chain-of-Thought reasoning process can be naturally integrated with KGs by translating a reasoning path on a KG into natural language. This flexibility allows us to easily test the model’s ability to follow instructions and make decisions based on the provided information. Similarly, In-Context Learning (ICL) [3] helps LLMs learn from demonstrative examples in the prompt to generate correct answers for the given question. This can also be naturally integrated with KGs. As a result, CoT and ICL enable flexible KG reasoning through natural language.
@@ 3.1 Knowledge Graph Definition
Let $KG=(E,R,L)$ denote a knowledge graph, where $E$ is the set of entities, $R$ is the set of relationships, and $L\subseteq E\times R\times E$ is the set of triples that are edges in the KG. Each triple $(e_{i},r,e_{i+1})\in L$ denotes that there exists a directed edge from entity $e_{i}$ to entity $e_{i+1}$ via the relationship $r$ [29].
@@ 3.1.1 Multi-hop Link Prediction
The task of multi-hop link prediction extends beyond simple link prediction between two nodes. It aims to identify missing connections over multiple relational steps within a knowledge graph. Specifically, given a sequence of observed triples that form a connected path $P_{obs}={(e_{1},r_{1},e_{2}),(e_{2},r_{2},e_{3}),\ldots,(e_{n-1},r_{n-1},e_{n})}\subseteq L$, where each triple $(e_{i},r_{i},e_{i+1})$ denotes an observed relation $r_{i}$ between entities $e_{i}$ and $e_{i+1}$. The objective is to predict whether a missing link $l_{miss}=(e_{1},?,e_{n})$ exists by answering True or False [19, 26].
@@ 3.1.2 Multi-hop Relation Prediction
The task of multi-hop relation prediction closely aligns with the concept of multi-hop link prediction, with a critical distinction in the question and output. Rather than determining the existence of a missing link $l_{miss}$, this task predicts the relationship. This change shifts the focus from a binary existence query to identifying the specific relationship that binds the entities [16].
@@ 3.2 Knowledge Prompt
In addition, we adopt the approach of utilizing one-shot ICL learning, specifically tailored to the FLAN-T5-Large, which is our smallest model. This is because, for models of this scale, the impact of utilizing one-shot ICL versus few-shot ICL on accuracy is minimal [3]. To maintain consistency across our experimental framework, we apply the same one-shot ICL methodology to all LLMs. This uniform approach ensures that our comparative analysis of the models’ performances is conducted under equivalent learning conditions. We listed all ICL examples in Appendix A.1.
@@ 3.3 KG-LLM Framework
Our complete KG-LLM Framework is illustrated in Figure 1. Initially, the KG is taken as input. Each node is iteratively assigned as the root, and depth-first search (DFS) is used to extract all possible paths. Duplicate paths are then removed, retaining only those with node counts ranging from 2 to 6. This range is based on the “six degrees of separation” theory [7], which states that any two individuals are, on average, connected through a chain of no more than six intermediaries. The node counts correspond to the number of hops: a single-hop is between two nodes, a two-hop involves three nodes, and so on. These paths are labeled as either positive (there is a connection between the first and last node) and negative (there is no connection) instances. We observed that negative instances outnumbered positive instances, so we randomly reduced the number of negative instances to achieve a balanced dataset. Finally, these paths are converted into KG-LLM and KG-LLM (ablation) knowledge prompts. During the fine-tuning phase, three distinct LLMs are utilized: Flan-T5-Large, Llama2-7B, and Gemma-7B. We added all the node IDs and relation IDs as special tokens to the vocabulary of these LLMs. Different fine-tuning techniques are applied for each model within our framework. A global fine-tuning strategy is employed on Flan-T5 to boost its performance. For Llama2 and Gemma, a 4-bit quantized LoRA (Low-Rank Adaptation) modification [10] is implemented. During the training process, we use the cross-entropy loss function $L$. It calculates the difference between the model’s predicted token probabilities and the actual token probabilities in the expected output sequence. In the following equation, $n$ represents the length of the expected output sequence, $x$ stands for the input instruction, and $y_{i}$ denotes the i-th token in the expected output sequence.
@@ 4.1 Experimental Setup
We conduct experiments over $four$ real-world datasets, WN18RR, NELL-995, FB15k-237 and YAGO3-10, which are constructed by the OpenKE library [9]. All datasets are commonly used for evaluating knowledge graph models in the field of knowledge representation learning. Statistics of the datasets are shown in Table 1.
@@
  • TransE [2] is a traditional distance model that represents relationships as translations in the embedding space.
  • Analogy [14] can effectively capture knowledge graph structures to improve link prediction.
  • CompleX [25] uses complex embeddings to represent both entities and relations, capturing asymmetric relationships.
  • DistMult [36] represents relations as diagonal matrices for simplicity and efficiency.
  • RESCAL [17] uses a tensor factorization method that captures rich interactions between entities and relations.
  • wsGAT [6] is a graph attention network that uses weighted self-attention mechanisms to perform various knowledge graph tasks.
  • ConGLR [12] leverages context-aware graph representation learning to enhance link prediction.
  • ConvRot [11] integrates convolutional networks and rotational embeddings to perform a variety of knowledge graph tasks.
@@
In evaluating the performance of models in multi-hop link prediction tasks, we utilized the Area Under the ROC Curve (AUC) metric [15] and the F1 score [20]. AUC measures the area under the Receiver Operating Characteristic (ROC) curve, which plots the true positive rate against the false positive rate at varying classification thresholds. The threshold is set at a 50% true positive rate and 50% false positive rate, as the number of positive and negative data points are equal in the testing case. A higher AUC value indicates a better ability of the model to differentiate between positive and negative examples. Similarly, the F1 score, ranging from 0 to 1, measures the balance between precision and recall, where higher values represent better performance. For the performance tables presented below, the best performance is indicated in bold, while the second-best performance is indicated with underline.
@@
(a) wsGAT
(b) ConGLR
(c) ConvRot
(d) Flan-T5 (Ablation)
(e) Llama 2 (Ablation)
(f) Gemma (Ablation)
(g) Flan-T5 (KGLLM)
(h) Llama 2 (KGLLM)
(i) Gemma (KGLLM)
Figure 4: Linear Relationship Between Complexity of Multi-Hop and Performance Score
@@ 4.2 Multi-hop Link Prediction without In-Context Learning
We also evaluate GNN, ablation, and KG-LLM framework models’ performance at each level of hop complexity in WN18RR and NELL-995 datasets. As shown in Figure 4, the performance of GNN and ablation models significantly declines as hop complexity increases. Upon closer examination, it is evident that as hop complexity grows, these models frequently respond with ’No’ for most questions, resulting in an F1 score close to 0 and an AUC score around 0.5. This performance is due to the increased complexities of multi-hop link prediction. Unlike the straightforward task of predicting a direct link between two nodes, models must consider all intermediate nodes to conclude, adding significant complexity and reducing their effectiveness. In contrast, the KG-LLM framework models effectively address this challenge, maintaining fair performance even at five-hops, except for the Flan-T5 model.
@@ 4.3 Multi-hop Link Prediction with In-Context Learning
(a) Ablation Framework
(b) KGLLM Framework
Figure 5: Multi-hop Relation Prediction Performance Comparison: The left graph shows model performance under the ablation framework, while the right graph shows model performance under the KGLLM framework. Blue bars represent testing without ICL, and red bars represent testing with ICL.
@@ 4.4 Multi-hop Relation Prediction without In-Context Learning
Our findings are presented in Table 5. We discovered that both frameworks showcased limited performance in this task without ICL. Notably, the KG-LLM framework exhibited marginally superior performance. Upon reviewing the predictive results, we observed that the model continues to provide ‘yes’ and ‘no’ answers for most questions, similar to the multi-hop link prediction task. For some questions, it outputs random responses.
@@ 4.5 Multi-hop Relation Prediction with In-Context Learning
We further explore the impact of incorporating ICL into the multi-hop relation prediction task. The ICL example is shown in the Appendix A.1. The results of red bars (with ICL) in Table 5 reveal a significant improvement in the generalization abilities of the models under both ablation and KG-LLM frameworks, in contrast to the blue bars (without ICL). In particular, the Llama 2 and Gemma models under the KG-LLM framework with ICL, achieved an accuracy exceeding 70% in the WN18RR datasets.

### v9

@@ Abstract
The task of multi-hop link prediction within knowledge graphs (KGs) stands as a challenge in the field of knowledge graph analysis, as it requires the model to reason through and understand all intermediate connections before making a prediction. In this paper, we introduce the Knowledge Graph Large Language Model (KG-LLM), a novel framework that leverages large language models (LLMs) for knowledge graph tasks. We first convert structured knowledge graph data into natural language and then use these natural language prompts to fine-tune LLMs to enhance multi-hop link prediction in KGs. By converting the KG to natural language prompts, our framework is designed to learn the latent representations of entities and their interrelations. To show the efficacy of the KG-LLM Framework, we fine-tune three leading LLMs within this framework, including Flan-T5, LLaMa2 and Gemma. Further, we explore the framework’s potential to provide LLMs with zero-shot capabilities for handling previously unseen prompts. Experimental results show that KG-LLM significantly improves the models’ generalization capabilities, leading to more accurate predictions in unfamiliar scenarios. Our code is available at https://github.com/rutgerswiselab/KG-LLM.
keywords:
Large Language Model, Knowledge Graph, Multi-Hop Link Prediction
editors: Vu Nguyen and Hsuan-Tien Lin
@@ 1 Introduction
Historically, approaches to solving tasks related to KGs can trace their origins from embedding-based methods to more recent advancements with LLMs (Wang et al., 2021). Initially, embedding-based methods played a crucial role, utilizing techniques to represent both entities and relations in a KG as low-dimensional vectors to address the link prediction task by preserving the structural and semantic integrity of the graph (Fan et al., 2014; Lin et al., 2015). As the field progressed, the integration of LLMs began to offer new paradigms, leveraging large amounts of data and advanced architectures to further enhance prediction capabilities and semantic understanding in KGs (Agarwal et al., 2020; Youn and Tagkopoulos, 2022; Yu et al., 2022; Tang et al., 2020). This transition shows a significant improvement from using purely mathematical representations to more context-aware methodologies that better understand the knowledge representation.
@@
To bridge these gaps, we propose the Knowledge Graph Large Language Model (KG-LLM), a novel approach to multi-hop link prediction. As illustrated in Figure 1, nodes in KGs are interconnected through specific relations. Initially, our framework takes input from the original KG dataset. After preprocessing, all paths in the KG will transform into chain-of-thought (CoT) prompts (Wei et al., 2022b; Jin et al., 2024), each includes a series of relational statements, which can be represented as {Node 1 (has relation_$x$ with) Node 2, Node 2 (has relation_$y$ with) Node 3, etc.}. The complexity of the multi-hop problem is determined by its path length and the number of nodes. Through instruction fine-tuning (IFT) (Wei et al., 2021) of three Large Language Models (LLMs): Flan-T5 (Wei et al., 2021), LlaMa2 (Touvron et al., 2023), and Gemma (Team et al., 2023), our framework is ready to enhance multi-hop link prediction performance during the testing phase. Moreover, by integrating in-context learning (ICL) (Xie et al., 2021), the model not only improves but also has the capacity to tackle unseen prompts, showcasing our method’s innovativeness in addressing multi-hop link prediction challenges.
@@ 2 Related Work
Recently, researchers have used Graph Neural Network (GNN) models to solve various graph-related tasks, significantly advancing the field. Among different GNN models, Graph Attention Networks (GATs) have gained attention for their ability to weigh the importance of neighboring nodes, with models like wsGAT (Grassia and Mangioni, 2022) demonstrating effectiveness in link prediction tasks. Additionally, Graph Convolutional Network (GCN)-based models have shown promising results; ConGLR (Lin et al., 2022) leverages context graphs and logical reasoning for improved inductive relation prediction, while ConvRot (Le et al., 2023) integrates relational rotation and convolutional techniques to enhance link prediction performance in knowledge graphs (KGs). While the aforementioned approaches have achieved significant success, multi-hop link prediction remains an unsolved challenge.
Other than GNN models, recent development of large language models (LLMs), such as BERT (Devlin et al., 2018), GPT (Peng et al., 2023), LLaMA (Touvron et al., 2023), Gemini (Team et al., 2023), and Flan-T5 (Wei et al., 2021) has also solved various KGs tasks, including link prediction. The text-to-text training approach makes LLMs particularly suitable for our generative multi-hop link prediction task. Recent studies, concurrent work such as GraphEdit (Guo et al., 2024), MuseGraph (Tan et al., 2024), and InstructGraph (Wang et al., 2024), have shown that natural language is effective for representing structural data for LLMs. Besides, training on large-scale data makes it possible for LLMs to generalize to unseen tasks or prompts that were not part of its training data (Wei et al., 2022a).
Another advantage of LLM-based generative modeling is the Chain-of-Thought (CoT) reasoning ability (Wei et al., 2022c). It provides the flexibility of modifying the instruction, options, and exemplars to allow structured generation and prediction. The Chain-of-Thought reasoning process can be naturally integrated with KGs by translating a reasoning path on a KG into natural language. This flexibility allows us to easily test the model’s ability to follow instructions and make decisions based on the provided information. Similarly, In-Context Learning (ICL) (Brown et al., 2020) helps LLMs learn from demonstrative examples in the prompt to generate correct answers for the given question. This can also be naturally integrated with KGs. As a result, CoT and ICL enable flexible KG reasoning through natural language.
@@ 3.1 Knowledge Graph Definition
Let $KG=(E,R,L)$ denote a knowledge graph, where $E$ is the set of entities, $R$ is the set of relationships, and $L\subseteq E\times R\times E$ is the set of triples that are edges in the KG. Each triple $(e_{i},r,e_{i+1})\in L$ denotes that there exists a directed edge from entity $e_{i}$ to entity $e_{i+1}$ via the relationship $r$ (Wang et al., 2017).
@@ 3.1.1 Multi-hop Link Prediction
The task of multi-hop link prediction extends beyond simple link prediction between two nodes. It aims to identify missing connections over multiple relational steps within a knowledge graph. Specifically, given a sequence of observed triples that form a connected path $P_{obs}={(e_{1},r_{1},e_{2}),(e_{2},r_{2},e_{3}),\ldots,(e_{n-1},r_{n-1},e_{n})}\subseteq L$, where each triple $(e_{i},r_{i},e_{i+1})$ denotes an observed relation $r_{i}$ between entities $e_{i}$ and $e_{i+1}$. The objective is to predict whether a missing link $l_{miss}=(e_{1},?,e_{n})$ exists by answering True or False (Ranganathan and Barbosa, 2022; Wan and Du, 2021).
@@ 3.1.2 Multi-hop Relation Prediction
The task of multi-hop relation prediction closely aligns with the concept of multi-hop link prediction, with a critical distinction in the question and output. Rather than determining the existence of a missing link $l_{miss}$, this task predicts the relationship. This change shifts the focus from a binary existence query to identifying the specific relationship that binds the entities (Nathani et al., 2019).
@@ 3.2 Knowledge Prompt
In addition, we adopt the approach of utilizing one-shot ICL learning, specifically tailored to the FLAN-T5-Large, which is our smallest model. This is because, for models of this scale, the impact of utilizing one-shot ICL versus few-shot ICL on accuracy is minimal (Brown et al., 2020). To maintain consistency across our experimental framework, we apply the same one-shot ICL methodology to all LLMs. This uniform approach ensures that our comparative analysis of the models’ performances is conducted under equivalent learning conditions. We listed all ICL examples in Appendix A.1.
@@ 3.3 KG-LLM Framework
Our complete KG-LLM Framework is illustrated in Figure 1. Initially, the KG is taken as input. Each node is iteratively assigned as the root, and depth-first search (DFS) is used to extract all possible paths. Duplicate paths are then removed, retaining only those with node counts ranging from 2 to 6. This range is based on the “six degrees of separation” theory (Guare, 2016), which states that any two individuals are, on average, connected through a chain of no more than six intermediaries. The node counts correspond to the number of hops: a single-hop is between two nodes, a two-hop involves three nodes, and so on. These paths are labeled as either positive (there is a connection between the first and last node) and negative (there is no connection) instances. We observed that negative instances outnumbered positive instances, so we randomly reduced the number of negative instances to achieve a balanced dataset. Finally, these paths are converted into KG-LLM and KG-LLM (ablation) knowledge prompts. During the fine-tuning phase, three distinct LLMs are utilized: Flan-T5-Large, LlaMa2-7B, and Gemma-7B. We added all the node IDs and relation IDs as special tokens to the vocabulary of these LLMs. Different fine-tuning techniques are applied for each model within our framework. A global fine-tuning strategy is employed on Flan-T5 to boost its performance. For LlaMa2 and Gemma, a 4-bit quantized LoRA (Low-Rank Adaptation) modification (Hu et al., 2021) is implemented. During the training process, we use the cross-entropy loss function $L$. It calculates the difference between the model’s predicted token probabilities and the actual token probabilities in the expected output sequence. In the following equation, $n$ represents the length of the expected output sequence, $x$ stands for the input instruction, and $y_{i}$ denotes the i-th token in the expected output sequence.
@@ 4.1 Experimental Setup
We conduct experiments over $four$ real-world datasets, WN18RR, NELL-995, FB15k-237 and YAGO3-10, which are constructed by the OpenKE library (Han et al., 2018). All datasets are commonly used for evaluating knowledge graph models in the field of knowledge representation learning. Statistics of the datasets are shown in Table 1.
@@
  • TransE (Bordes et al., 2013) is a traditional distance model that represents relationships as translations in the embedding space.
  • Analogy (Liu et al., 2017) can effectively capture knowledge graph structures to improve link prediction.
  • CompleX (Trouillon et al., 2016) uses complex embeddings to represent both entities and relations, capturing asymmetric relationships.
  • DistMult (Yang et al., 2014) represents relations as diagonal matrices for simplicity and efficiency.
  • RESCAL (Nickel et al., 2016) uses a tensor factorization method that captures rich interactions between entities and relations.
  • wsGAT (Grassia and Mangioni, 2022) is a graph attention network that uses weighted self-attention mechanisms to perform various knowledge graph tasks.
  • ConGLR (Lin et al., 2022) leverages context-aware graph representation learning to enhance link prediction.
  • ConvRot (Le et al., 2023) integrates convolutional networks and rotational embeddings to perform a variety of knowledge graph tasks.
@@
In evaluating the performance of models in multi-hop link prediction tasks, we utilized the Area Under the ROC Curve (AUC) metric (Lobo et al., 2008) and the F1 score (Sai et al., 2022). AUC measures the area under the Receiver Operating Characteristic (ROC) curve, which plots the true positive rate against the false positive rate at varying classification thresholds. The threshold is set at a 50% true positive rate and 50% false positive rate, as the number of positive and negative data points are equal in the testing case. A higher AUC value indicates a better ability of the model to differentiate between positive and negative examples. Similarly, the F1 score, ranging from 0 to 1, measures the balance between precision and recall, where higher values represent better performance. For the performance tables presented below, the best performance is indicated in bold, while the second-best performance is indicated with underline.
@@ 4.2 Multi-hop Link Prediction without In-Context Learning
We also evaluate GNN, ablation, and KG-LLM framework models’ performance at each level of hop complexity in WN18RR and NELL-995 datasets. As shown in Figure LABEL:linear_relation, the performance of GNN and ablation models significantly declines as hop complexity increases. Upon closer examination, it is evident that as hop complexity grows, these models frequently respond with ’No’ for most questions, resulting in an F1 score close to 0 and an AUC score around 0.5. This performance is due to the increased complexities of multi-hop link prediction. Unlike the straightforward task of predicting a direct link between two nodes, models must consider all intermediate nodes to conclude, adding significant complexity and reducing their effectiveness. In contrast, the KG-LLM framework models effectively address this challenge, maintaining fair performance even at five-hops, except for the Flan-T5 model.
@@ 4.4 Multi-hop Relation Prediction without In-Context Learning
Our findings are presented in Table LABEL:relation_prediction. We discovered that both frameworks showcased limited performance in this task without ICL. Notably, the KG-LLM framework exhibited marginally superior performance. Upon reviewing the predictive results, we observed that the model continues to provide ‘yes’ and ‘no’ answers for most questions, similar to the multi-hop link prediction task. For some questions, it outputs random responses.
@@ 4.5 Multi-hop Relation Prediction with In-Context Learning
We further explore the impact of incorporating ICL into the multi-hop relation prediction task. The ICL example is shown in the Appendix A.1. The results of red bars (with ICL) in Table LABEL:relation_prediction reveal a significant improvement in the generalization abilities of the models under both ablation and KG-LLM frameworks, in contrast to the blue bars (without ICL). In particular, the LlaMa2 and Gemma models under the KG-LLM framework with ICL, achieved an accuracy exceeding 70% in the WN18RR datasets.
@@ A.1.4 Multi-hop Relation Prediction ICL Example in KG-LLM Framework
A.2 Dataset 2
A.2.1 ablation Framework One-shot In-Context Learning for Multi-hop Link Prediction
### Context:
node_11809 has relation_9 with node_12218. node_12218 has relation_1 with node_9431. Is node 11809 connnected with node 9431?
Answer:
The answer is yes.
A.2.2 KG-LLM Framework One-shot In-Context Learning for Multi-hop Link Prediction
### Context:
node_11809 has relation_9 with node_12218. node_12218 has relation_1 with node_9431. Is node 11809 connnected with node 9431?
Answer:
node_11809 has relation_9 with node_12218, means node_11809 _verb_group node_12218. node_12218 has relation_1 with node_9431, means node_12218 _derivationally_related_form node_9431. So node 11809 _derivationally_related_form node 9431. The answer is yes.
A.2.3 ablation Framework One-shot In-Context Learning for Multi-hop Relation Prediction
### Context:
node_11809 has relation_9 with node_12218. node_12218 has relation_1 with node_9431. What is the relationship between node_11809 and node_9431?
Answer:
The relationship between the first node and the last node is relation_1.
A.2.4 KG-LLM Framework One-shot In-Context Learning for Multi-hop Relation Prediction
### Context:
node_11809 has relation_9 with node_12218. node_12218 has relation_1 with node_9431. What is the relationship between node_11809 and node_9431?
Answer:
node_11809 has relation_9 with node_12218, means node_11809 _verb_group node_12218. node_12218 has relation_1 with node_9431, means node_12218 _derivationally_related_form node_9431. So node 11809 _derivationally_related_form node 9431. The relationship between the first node and the last node is _derivationally_related_form.